OpenStudy (yvonneam0304):

Fiona is serving iced tea and lemonade at a picnic. She has only 44 cups in which to serve the drinks. If x represents the number of glasses of iced tea and y represents the number of glasses of lemonade, which equation represents the number of drinks she can serve

OpenStudy (renee9906):

X+y=44 (number of glasses of iced tea)+(number of glasses of lemonade)=44 drinks total x + y = 44 so, x+y=44 is your equation. OR, if you want it in the formula of y=mx+b, then- -x -x y=-x+44

likeabossssssss (likeabossssssss):

welcome to os